Sample size is n A. {lto/> 1.895} B. {lto?> 1.833) C. ltol > 1.415) D. {Ito/> 1.383) 9. 9. what is the rejection region at the ?-0.10 level? 10. Which of the following statements is NOT true? A. B. C. An important use of control charts is to check if the process average remains stable over time. The R chart is used to monitor changes in the process variance ?2 If a process is stable or in control, the probability that the sample mean X will fall between and + 3?An is very small. -3?An D. The u chart is used to monitor the average number of defectives (e.g. dents, scratches) found in each unit.
Answer #1

9.

n =9

So,

ndf = 9 - 1 = 8

\alpha = 0.10

From Table,critical value of t = \pm 1.8595

Rejection region:

Correct option:

A.    \begin{vmatrix} t_{0} \end{vmatrix}>1.895

10.

Option:

D.

EXPLANATION:

u chart is used only for defectives and not defects.